OpenStudy (anonymous):

how to prove that the limit e->+0 of integral (-inf,+inf) of (f(x)*exp(-x^2/4e)/(2pi*sqrt(e))) equals f(0)

OpenStudy (anonymous):

i have no idea, but i am going to try and write it \[\large \lim_{\epsilon\to 0^+}\int_{-\infty}^{\infty}f(x)\frac{\exp(\frac{x^2}{4\epsilon})}{2\pi\sqrt{\epsilon}}=f(0)\]

OpenStudy (anonymous):

Thank you, satellite73, the way I wrote is unreadable

OpenStudy (anonymous):

oh, and f(x) has derivatives of every N, all of them continuous, and there is R, such as f(x)=0 for any x: abs(x)>R

ganeshie8 (ganeshie8):

can i change epsilon/e to t ? \[\large \lim_{t\to 0^+}\int_{-\infty}^{\infty}f(x)\frac{\exp(\color{red}{-}\frac{x^2}{4t})}{2\pi\sqrt{t~}}=f(0) \]

OpenStudy (anonymous):

of course

ganeshie8 (ganeshie8):

does dominated convergence thm apply here ?

OpenStudy (anonymous):

well, we are not supposed to use it, as we did not have it in our lectures

OpenStudy (anonymous):

I have just looked through it on wiki, I think I can use it

OpenStudy (anonymous):

um, ganeshie8, I see you have received a medal, that probably means that your post is practically the answer, but I still don't get it. Could you please explain it to me? :)

ganeshie8 (ganeshie8):

Hey no, I am kind of stuck. Let me tag few @eliassaab @Alchemista

OpenStudy (anonymous):

http://math.stackexchange.com/questions/55137/is-this-a-delta-function-and-delta-as-limit-of-gaussian I have found the solution, question is closed